Question 1)

Refer to the Karla Inc. information below. Karla Inc. manufactures tents under various brand names. The company sells most of its hammocks in the second quarter of each year. Their production budget for the second quarter shows the following number of hammocks needs to be produced:

April 7000 units

May 15 000 units

June 25 000 units
Each unit requires 25 feet of cotton rope cord which costs $0.60 per foot. The company has determined that it needs 20 per cent of next month’s raw material needs on hand at the end of each month. In addition, each hammock requires 55 minutes of direct labour for assembly and inspection at a cost of $0.45 per minute. The company currently applies manufacturing overhead to production at the rate of $9.50 per direct labour hour. The total cost of direct labour and manufacturing overhead for the month of May is:


a) $201 775
b) $501 875
c) $635 201
d) $501 775
Answer #1

Answer : b) $ 501875

Number of units to be produced for the month of May = 15000 units

Labor time required for each hammock = 55 minutes

Total direct labor hours needed for the month of may = 15000 * 55 = 825000 minutes

Cost of direct labor for assembly and inspection = $.45 per minute

total cost of direct labor for May = 825000 * .45 = $ 371250

Manufacturing overhead applying rate to production = $ 9.5 per direct labor hour

total direct labor hour required for may= 825000 minutes/ 60 = 13750 hours

therefore total manufacturing overhead for the month May = 13750 * 9.5 = $ 130625

Total cost of direct labor hour and manufacturing overhead for the month of May

= 371250 + 130625 = $ 501875